OpenStudy (anonymous):

(13a+1)-(5a+13)=-6 find the value of a

OpenStudy (anonymous):

what choices is for the question?

OpenStudy (anonymous):

First step is to distribute that negative in the middle through the (5a + 13) (13a +1)-(5a +13)= -6 (13a + 1) - 5a - 13 = -6 With me so far?

OpenStudy (anonymous):

Then we can remove the parentheses from the (13a + 1) To get to 13a + 1 -5a - 13 = -6 Now combine like terms.

OpenStudy (anonymous):

Combineing like terms, 13a - 5a = 8a 1 - 13 = -12 We get 8a - 12 = -6 Add 12 to both sides... 8a = 6 Can you solve for a now?

OpenStudy (anonymous):

a=3/4

OpenStudy (anonymous):

Now check you answer by plugging in (3/4) into where the a's are and see if it comes out to be a true statement. This is how you will know if you are correct or not.

OpenStudy (anonymous):

[13(3/4)+1]-[5(3/4)+13]=-6 10.75 - 16.75 = -6 -6 = -6 True! So a= 3/4 is correct! =0)

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

Np man, any time.